Show pageDiscussionOld revisionsBacklinksBack to top This page is read only. You can view the source, but not change it. Ask your administrator if you think this is wrong. ======20231101========== =====Time===== 1800-0200(+1) =====Attendance===== Onkar, Roy =====Spaces Occupied===== PHYS 401, 433C, Dome =====Goals===== * Clean-up data on DC/External SSD/gDrive * Take frames for tracking * Take frames for seeing * Run through process for determining limiting magnitude * <del> Attach PICSARR to scope </del> @1820, then nvm scrub @1907, then nvm still scrubbed @1930 =====Frame Table===== ^ Time Observed ^ Source Name ^ Filter ^ Int Time [s] ^ Mag (filter) ^ RA [h] ^ DEC [deg] ^ Source Type ^Scope ^ Purpose ^ Notes^ | 20231101T1957| Vega (HD 172167) | V | 0.864 | 0(V) | 18 | 38 | Star |Scope| Pointing | None | |20231101T2057|M15|R|30|6.30(V)|21|12|Cluster|Main|Tracking | 15.042w/o-ag| |20231101T2131|M15|R|30|6.30(V)|21|12|Cluster|Finder|Tracking |15.042w/o-ag| |20231101T2200|M15|R|30|6.30(V)|21|12|Cluster|Main|Tracking|15.042w/ag| |20231101T2213|M15|R|30|6.30(V)|21|12|Cluster|Main|Tracking|15.009w/ag| |20231101T22|M15|R|30|6.30(V)|21|12|Cluster|Main|Tracking|15.011w/ag| |20231101T2300|HD 222603|V|4.5|4.5(V)|23.75|2|Star|Main|Seeing|15.011 w/ag| |20231101T2330|HD 41927|V|10|5.36(V)|6|65|Star|Main|Seeing|15.011 w/ag| |20231101T2337|$\phi~Cas$ (HD 7927)|V|10|4.95(V)|1|58|Star|Main|Seeing/Mag|15.011 w/ag| =====Log===== - Data Clean-Up - Confirmed all D-Drive data on DC was synced to GoogleDrive. - Removed all D-Drive data on DC - Renamed D-Drive HDD-Drive - Normal Start-Up Procedure - Pointed at Vega - Moved to M15 to collect tracking frames - Set RA Tracking Rate to 15.042"/s and collected 10 frames with main scope without autoguiding - Set RA Tracking Rate to 15.042"/s and collected 10 frames with finder without autoguiding - Set RA Tracking Rate to 15.042"/s and collected 10 frames with main scope with autoguiding - Set RA Tracking Rate to 15.009"/s and collected 10 frames with main scope with autoguiding - Set RA Tracking Rate to 15.011"/s and collected 10 frames with main scope with autoguiding - Turned off all security camera IR lamps - Moved to HD 222603 and took 3x light frames TR=15.011"/s - Moved to HD 41927 and took 3x light frames TR=15.011"/s - Moved to ($\phi~Cas$)HD 7927 and stepped through smart histogram tool outlined by Eileen TR=15.011"/s - Recorded Smart Histogram Results for each filter - Took 10 frames in each filter with recommended settings from Smart Histogram - Took 10x dark frames with each integration time mentioned above =====ToDo===== =====Notes===== - North Mirror Door Assembly was misbehaving (not closing when asked by DFM TCS) - Onkar reported high-pitched noise from motor assembly when closing - We were able to force all doors closed - Then they opened and closed successfully, with North slightly slower - Onkar will investigate tomorrow (Thurs) - TS6 did not take the scope to zenith -- it was about $10^\circ$ off of zenith - We were able to get DFM TCS to move the scope to zenith - Telescope > Movement > Offset/Zenith > Set Zenith Position (hit Apply) > then "Start Slew" =====Weather===== =====Data===== - Darks stored in SharpCap Captures in gDrive: 2023-11-01 with "dark_#". Where #/10 is the integration time in s - FLATS and Limiting Magnitude frames are in Projects > ASI432MM_bigpix > 20231102 along with all light frames